# Hierarchical Relation-augmented Representation Generalization for Few-shot Action Recognition

### v1

@@ Abstract.
Few-shot action recognition (FSAR) aims to recognize novel action categories with few exemplars. Existing methods typically learn frame-level representations independently for each video by designing various inter-frame temporal modeling strategies. However, they neglect explicit relation modeling between videos and tasks, thus failing to capture shared temporal patterns across videos and reuse temporal knowledge from historical tasks. In light of this, we propose HR2G-shot, a Hierarchical Relation-augmented Representation Generalization framework for FSAR, which unifies three types of relation modeling (inter-frame, inter-video, and inter-task) to learn task-specific temporal patterns from a holistic view. In addition to conducting inter-frame temporal interactions, we further devise two components to respectively explore inter-video and inter-task relationships: i) Inter-video Semantic Correlation (ISC) performs cross-video frame-level interactions in a fine-grained manner, thereby capturing task-specific query features and learning intra- and inter-class temporal correlations among support features; ii) Inter-task Knowledge Transfer (IKT) retrieves and aggregates relevant temporal knowledge from the bank, which stores diverse temporal patterns from historical tasks. Extensive experiments on five benchmarks show that HR2G-shot outperforms current top-leading FSAR methods. Full code will be released.
copyright: none
1. Introduction
@@
During the last few years, action recognition (Feichtenhofer et al., 2019; Wang et al., 2021a; Xing et al., 2023a; Huang et al., 2023, 2024a; Cao et al., 2024; Li et al., 2024) has witnessed remarkable progress with the advances in deep learning. However, their strong performance heavily relies on a large amount of labeled training examples, which restricts the model’s scalability and generalizability in data scarcity scenarios. In contrast, human beings can easily learn new visual concepts with only a few supervisions. To enable the machine to acquire such ability, many studies have shifted their attention towards few-shot action recognition (FSAR) (Zhu and Yang, 2018; Cao et al., 2020; Wang et al., 2024c), which aims to learn novel (unseen) action classes using only a few annotated video samples after training on a set of base (seen) classes with abundant samples.
@@
Current top-leading FSAR solutions typically adopt the metric-based meta-learning paradigm, wherein the model first maps the query videos and support videos into discriminative feature space, and then performs query-support video matching based on pre-defined or learned distance metrics. To achieve this, a prevalent of subsequent efforts delve into network designs for temporal representation learning by, e.g., temporal attention operations (Wang et al., 2022), elaborated spatio-temporal (Thatipelli et al., 2022) or temporal-channel interaction (Xing et al., 2023c), multi-layer feature fusion (Xing et al., 2023b), and auxiliary modal information (e.g., depth (Fu et al., 2020) and motion (Wang et al., 2023a)). Apart from investigating temporal representation learning, to perform proper temporal feature alignment when comparing query and support videos, recent approaches devise various temporal matching strategies, e.g., frame-level feature matching (Cao et al., 2020; Wang et al., 2024c), tuple-level feature matching (Perrett et al., 2021), frame-to-segment feature matching (Wu et al., 2022), or multi-level feature matching (Zheng et al., 2022).
Though achieving remarkable performance, these methods suffer from two limitations: First, they independently learn frame-level representations for each video, neglecting the relation modeling and holistic semantic patterns among video samples in the current task. Exploring inter-video relationships can not only learn subtle semantic discrepancies among different videos, but also capture shared similar temporal patterns (e.g., action styles and speed) from diverse video samples, particularly under data-limited conditions. Although some preliminary attempts (Xing et al., 2023c; Wang et al., 2022) have been made towards this, they capture sample-level correlations at the video level (global feature for each video), ignoring fine-grained temporal details in videos. Thus, these methods fail to capture shared temporal patterns across videos in each task.
@@
Figure 1. Our main idea. Previous FSAR works only rely on inter-frame relation modeling to learn video representations, ignoring the relations between videos and tasks. In contrast, we unify three types of relation modeling (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to capture task-specific temporal cues.
Second, existing methods directly learn meaningful video representation from one single task, lacking explicit modeling of task-relatedness (i.e., some tasks share similar action patterns or structures). Thus, they fail to reuse experience from historical tasks and excavate task-shared temporal knowledge, causing sub-optimal transferability and generalization on new tasks with only a few samples. In light of the above, our goal is to unify three types of relation modeling (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to yield task-specific temporal features for each video via transferring knowledge at both video-level and task-level (as shown in Fig. 1).
@@
In this vein, we develop a novel hierarchical framework for FSAR, namely HR2G-shot, which learns task-specific temporal cues from multiple perspectives (i.e., inter-frame, inter-video, inter-task) in a progressive manner. In addition to performing inter-frame temporal interactions, we further explicitly learn inter-video and inter-task relationships, so as to strengthen temporal pattern understanding for each video from a holistic view. Thus, we propose two feasible modules: i) Inter-video Semantic Correlation (ISC) aims to precisely capture inter-video temporal relationships within each task. Specifically, we perform cross-video frame-level interactions in a fine-grained manner to gain task-specific temporal features and reduce considerable computational costs compared with inter-video dense interaction (see Table 3). Moreover, to respect the difference between support and query videos, we customize the masked interaction strategy for support-support and query-support relations (i.e., support and query videos only aggregate semantic information from support videos). By this means, our model could interactively transfer temporal knowledge between videos to yield task-specific query features and learn intra- and inter-class correlation among support features. ii) Inter-task Knowledge Transfer (IKT) is devised to learn transferable temporal knowledge between tasks by exploring inter-task relationships. Concretely, we develop a temporal knowledge bank to store diverse temporal patterns from historical tasks. Then we make use of temporal prototypes, which summarize frame-level representations of each video, to retrieve and aggregate temporal knowledge related to new tasks from the bank. Integrated with such knowledge, our method yields more meaningful video representations, thus quickly adapting to new tasks with only a few samples. We conduct extensive experiments on five gold-standard datasets, and the results demonstrate that our HR2G-shot outperforms existing top-leading FSAR methods by a large margin.
@@
  • We provide a hierarchical relation-augmented framework for FSAR, which unifies three types of relation modeling to learn task-specific temporal patterns.
@@
  • To explore inter-video temporal relationships within each task, we propose Inter-video Semantic Correlation that conducts cross-video frame-level interactions in a fine-grained manner and customizes masked interaction strategy for support-support and query-support relations.
@@
  • To learn transferable temporal knowledge between tasks, we design Inter-task Knowledge Transfer that retrieves and aggregates relevant temporal knowledge from the temporal knowledge bank, which stores diverse temporal patterns from historical tasks.
2. Related Work
@@
Few-shot Image Classification. The goal of few-shot image classification (Fei-Fei et al., 2006) is to recognize unseen categories with limited annotated samples. Existing few-shot solutions can be broadly categorized into three groups: i) Augmentation-based methods typically generate auxiliary data samples by feature augmentation (Chen et al., 2019a, b; Albalak et al., 2024) or generative models (Li et al., 2020; Zhang et al., 2018), to increase the diversity of the support set; ii) Optimization-based methods (Jamal and Qi, 2019; Sun and Gao, 2024; Rusu et al., 2019; Rajeswaran et al., 2019; Finn et al., 2017) learn good initial model parameters, such that updating the initial parameters via a few gradient steps could adapt well to new tasks; iii) Metric-based methods (Yoon et al., 2019; Yang et al., 2024; Zhu and Koniusz, 2023; Cheng et al., 2023) (e.g., ProtoNet (Snell et al., 2017) and RelationNet (Sung et al., 2018)) first learn a common feature space for all classes, and then employ pre-defined or learned distance metrics to compare query and support samples.
@@
Our work shares a similar spirit of metric-based methods (Snell et al., 2017; Vinyals et al., 2016), whereas we address more complex and challenging few-shot action recognition that requires long-range relation modeling between spatio-temporal locations. With respect to this, we further jointly learn inter-video and inter-task relationships to learn distinct temporal features. By unifying such three types of relation modeling under our framework, we can learn task-specific temporal cues for each video from a holistic view.
3.3)3.4)$\bm{X}_{\mathrm{s}},\bm{X}_{\mathrm{q}}$$\bm{F}_{\mathrm{s}},\bm{F}_{\mathrm{q}}$$\tilde{\bm{F}}_{\mathrm{s}},\tilde{\bm{F}}_{\mathrm{q}}$(Eq. 7)
Figure 2. The overview of HR2G-shot. (a) HR2G-shot unifies three types of relation modeling (i.e., inter-frame, inter-video, and inter-task) to learn discriminative temporal features. (b) Inter-video Semantic Correlation (ISC) conducts fine-grained cross-video interactions to learn inter-video relationships. (c) To explore inter-task relationships, we retrives and aggregates temporal knowledge from the bank, which maintains diverse temporal patterns from historical tasks.
@@
Few-shot Action Recognition (FSAR). FSAR is gaining much attention due to its practical value in reducing manual annotations at a large scale required by traditional action recognition. Compared with few-shot image classification, FSAR is more challenging due to the rich spatial-temporal information in videos. Towards FSAR, most existing solutions (Wang et al., 2024a; Wu et al., 2024; Liu et al., 2022; Tang et al., 2023) belong to the meta-learning paradigm (Snell et al., 2017; Vinyals et al., 2016), wherein the model learns discriminative temporal features from base action categories, and generalizes learned temporal representations on novel action categories. Based on the meta-learning policy, a main group of efforts focus on temporal representation learning. Early works (Zhu and Yang, 2018, 2020) directly aggregate frame-level features to learn a global representation for each video. Though straightforward, these methods neglect complex temporal cues within one video, leading to suboptimal performance. To address this limitation, recent methods attempt to model temporal relations by, e.g., learning long-range temporal dependencies (Thatipelli et al., 2022; Wang et al., 2022), detailed cross-frame patch-level interactions (Zhang et al., 2023), exploiting low-level spatial features via feature architecture search (Xing et al., 2023b), or multi-modal feature fusion (Wanyan et al., 2023; Fu et al., 2020; Wang et al., 2021b; Huang et al., 2024c). Apart from investigating video feature representation, another line of work turns to designing various feature matching strategies, e.g., frame-level feature alignment (Wang et al., 2022), tuple-level feature alignment (Perrett et al., 2021), and even multi-level feature alignment (Zheng et al., 2022; Wu et al., 2024; Huang et al., 2024b). For example, OTAM (Cao et al., 2020) performs frame-level feature alignment between videos with the DTW algorithm, which takes the temporal ordering characteristics of videos into account. HyRSM (Wang et al., 2022) designs a novel bidirectional Mean Hausdorff metric to obtain temporal matching scores between support and query videos from the set matching perspective.
@@
Despite promising results have been achieved, they typically learn temporal representations independently for each video, ignoring the relation modeling between video samples and tasks. As a result, they fail to gain a holistic semantic pattern understanding from multi-level perspectives. Though a few works (e.g., GgHM (Xing et al., 2023c) and HyRSM (Wang et al., 2022)) attempt to learn task-specific embeddings by considering inter-task relationships, they only focus on the correlations between samples at the video level, overlooking fine-grained temporal interactions among videos. In contrast, our HR2G-shot i) learns inter-video relationships at fine-grained frame level, and ii) gather temporal knowledge from historical tasks to achieve inter-task knowledge transfer, thereby capturing task-specific temporal features for each video and ensuring a sufficient understanding of holistic semantic patterns for each task.
3. Method
3.1. Problem Formulation
@@
The objective of few-shot action recognition (FSAR) is to classify novel action categories with limited samples per class. Under the few-shot setting, the model generalizes learned knowledge on training set $\mathcal{D}_{base}$ to novel testing set $\mathcal{D}_{novel}$, where base classes $\mathcal{C}_{base}$ and novel classes $\mathcal{C}_{novel}$ are disjoint. Consistent with previous works (Cao et al., 2020; Wang et al., 2022; Tang et al., 2024), we formulate FSAR problem with episodic training and testing . Specifically, each $N$-way $K$-shot episode task involves sampled $K$ labeled videos from each of $N$ different action classes as the support set, and a portion of the remaining samples from $N$ classes as the query set. During inference, we randomly sample multiple episodic tasks from $\mathcal{D}_{novel}$ and present the average results over these tasks, in order to comprehensively evaluate the performance of the few-shot model.
3.2. Algorithm Overview
@@
We introduce HR2G-shot, which hierarchically models multi-level relations (i.e., inter-frame, inter-video, and inter-task) to capture task-specific temporal patterns from a holistic view for FSAR. The overall architecture of HR2G-shot is shown in Fig. 2. Given an $N$-way $K$-shot episode task with $NK$ support videos and $L$ query videos, we employ CLIP visual encoder (Radford et al., 2021) to get frame-level features for support and query videos, i.e., $\bm{X}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{X}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$, where $T$ is the length of sampled frames in each video. Note that we adopt Parameter-Efficient Fine-Tuning (PEFT) strategy to fine-tune the visual encoder with minimal trainable parameters as in (Xing et al., 2024; Qu et al., 2025; Yang et al., 2023). Next, following previous works (Wang et al., 2024b), we concatenate frame-level features and corresponding textual features along the temporal dimension and then feed into temporal Transformer (Wang et al., 2024b, 2023a) (i.e., Inter-frame Temporal Modeling). In this way, we perform inter-frame interactions and obtain enhanced support and query features. To compensate frame-level relation modeling from other semantic perspectives, we further design two feasible modules: i) Inter-video Semantic Correlation (ISC) (§3.3) learns inter-video temporal relationships within each task by fine-grained cross-video interactions; ii) Inter-task Knowledge Transfer (IKT) (§3.4) aggregates useful temporal knowledge from historical tasks to learn inter-task relationships. Finally, the obtained task-specific support and query features, i.e., $\bar{\bm{F}}_{\mathrm{s}}$ and $\bar{\bm{F}}_{\mathrm{q}}$, are fed into temporal matching metrics to get the class predictions.
3.3. Inter-video Semantic Correlation

### v3

@@ Abstract
Few-shot action recognition (FSAR) aims to recognize novel action categories with limited exemplars. Existing methods typically learn frame-level representations independently for each video by designing various inter-frame temporal modeling strategies. However, these methods suffer from “cognitive limitation” regarding information isolation, treating each video and task as an independent entity. This information isolation contradicts human cognitive mechanisms, which dynamically associate the current observation with concurrent examples to distill common temporal patterns (inter-video correlation) and recall relevant motion priors from long-term memory (inter-task knowledge transfer). In light of this, we propose HR2G-shot, a Hierarchical Relation-augmented Representation Generalization framework for FSAR, which mimics the human cognitive process by unifying three types of relation modeling (inter-frame, inter-video, and inter-task) to learn task-specific temporal patterns from a holistic view. In addition to conducting inter-frame temporal interactions, we further devise two cognitively-inspired components to respectively explore inter-video and inter-task relationships: i) Inter-video Semantic Correlation (ISC) performs cross-video frame-level interactions in a fine-grained manner, thereby capturing task-specific query features and learning intra- and inter-class temporal correlations among support features; ii) Inter-task Knowledge Transfer (IKT) retrieves and aggregates relevant temporal knowledge from the bank, which stores diverse temporal patterns from historical tasks. Extensive experiments on five benchmarks show that HR2G-shot outperforms current top-leading FSAR methods.
keywords:
Action Recognition; Video Understanding; Few-shot Action Recognition; Few-shot Learning
1 Introduction
@@
During the last few years, action recognition [1, 2, 3, 4, 5, 6, 7, 8] has witnessed remarkable progress with the advances in deep learning. However, their strong performance heavily relies on a large amount of labeled training examples, which restricts the model’s scalability and generalizability in data scarcity scenarios. In contrast, human beings can easily learn new visual concepts with only a few supervisions. To enable the machine to acquire such ability, many studies have shifted their attention towards few-shot action recognition (FSAR) [9, 10, 11, 12], which aims to learn novel (unseen) action classes using only a few annotated video samples after training on a set of base (seen) classes with abundant samples.
@@
Current top-leading FSAR solutions typically adopt the metric-based meta-learning paradigm, wherein the model first maps the query videos and support videos into a discriminative feature space, and then performs query-support video matching based on pre-defined or learned distance metrics. To achieve this, a prevalent number of subsequent efforts delve into network designs for temporal representation learning by, e.g., temporal attention operations [13], elaborated spatio-temporal [14] or temporal-channel interaction [15], multi-layer feature fusion [16], and auxiliary modal information (e.g., depth [17] and motion [18]). Apart from investigating temporal representation learning, to perform proper temporal feature alignment when comparing query and support videos, recent approaches devise various temporal matching strategies, e.g., frame-level feature matching [10, 11], tuple-level feature matching [19], frame-to-segment feature matching [20], or multi-level feature matching [21].
Despite their performance, we argue that these methods suffer from “cognitive limitation” regarding information isolation (Fig. 1a). First, existing methods directly learn meaningful video representations from one single task, lacking an explicit mechanism for longitudinal knowledge accumulation (inter-task relations). Consequently, they fail to distill universal temporal patterns that are shared across diverse actions. Actually, novel actions are often variations or compositions of previously learned motion primitives. Ignoring this inter-task relatedness prevents the model from “warm-starting” with prior experience, causing sub-optimal transferability and generalization on new tasks with only a few samples.
@@
Figure 1: Our main idea. (a): Previous FSAR works only rely on inter-frame relation modeling to learn video representations, ignoring the relations between videos and tasks. (b): In contrast, we unify three types of relation modeling (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to capture task-specific temporal cues.
Second, they independently learn frame-level representations for each video, neglecting the collective semantics (inter-video relations) in the current task. Without cross-referencing multiple examples, models struggle to distinguish task-relevant motion patterns from instance-specific noise (e.g., complex backgrounds or camera motion). Exploring inter-video relationships allows the model to reach a “consensus” on the common action essence (e.g., action styles and speed) while filtering out non-shared discrepancies. Although some preliminary attempts [15, 13] have been made towards this, they capture sample-level correlations at the video level (global feature for each video), ignoring fine-grained temporal details in videos. Thus, these methods fail to capture how specific motion segments align across different samples.
From a cognitive perspective, these limitations stem from a fundamental deviation from human intelligence. Human capability is rooted in a holistic cognitive mechanism [22]: when humans learn a new action, they do not view events in isolation. Instead, they dynamically associate the current observation with concurrent examples to distill common patterns (contextualization) [23] and recall relevant motion priors from long-term memory (knowledge transfer) [24]. Without such a mechanism, existing models struggle to capture the intrinsic structure of actions from a few examples, creating a significant gap in sample efficiency compared to human learners. In light of the above, our goal is to shift the paradigm from isolated instance learning to hierarchical relational learning (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to yield task-augmented temporal features for each video via transferring knowledge at both video-level and task-level (as shown in Fig. 1b).
@@
In this vein, we develop a novel hierarchical framework for FSAR, namely HR2G-shot, which mimics the human cognitive process by progressively modeling temporal cues across three perspectives: micro-level (inter-frame), meso-level (inter-video), and macro-level (inter-task). By unifying these hierarchies, we enable the model to not only parse internal motion details (i.e., inter-frame temporal interactions) but also make use of external relationships for holistic understanding. To achieve this, we propose two cognitively-inspired modules: i) Inter-video Semantic Correlation (ISC) aims to precisely capture meso-level contextualization within each task. Specifically, we perform cross-video frame-level interactions in a fine-grained manner to gain task-specific temporal features and reduce considerable computational costs compared with inter-video dense interaction (see Table 5). Moreover, to respect the difference between support and query videos, we customize the masked interaction strategy for support-support and query-support relations (i.e., support and query videos only aggregate semantic information from support videos). This asymmetric design ensures that query videos actively aggregate semantic cues from support anchors, while support videos reinforce their intra-class compactness, thereby yielding robust task-specific representations. ii) Inter-task Knowledge Transfer (IKT) is devised to realize macro-level knowledge transfer between tasks by exploring inter-task relationships. Concretely, we develop a temporal knowledge bank that serves as the model’s long-term memory, storing diverse temporal patterns from historical tasks. Then we make use of temporal prototypes, which summarize frame-level representations of each video, to retrieve and aggregate temporal knowledge related to new tasks from the bank. Integrated with such knowledge, our method yields more meaningful video representations, thus quickly adapting to new tasks with only a few samples. We conduct extensive experiments on five gold-standard datasets, and the results demonstrate that our HR2G-shot outperforms existing top-leading FSAR methods by a large margin.
@@
  • We provide a hierarchical relation-augmented framework for FSAR, which mimic the human cognitive process by progressively modeling temporal cues across three perspectives: micro-level (inter-frame), meso-level (inter-video), and macro-level (inter-task).
@@
  • To capture collective semantics within each task, we propose Inter-video Semantic Correlation module that conducts cross-video frame-level interactions in a fine-grained manner and customizes masked interaction strategy for support-support and query-support relations.
@@
  • To enable longitudinal knowledge accumulation, we design Inter-task Knowledge Transfer module that retrieves and aggregates relevant temporal knowledge from the temporal knowledge bank, which stores diverse temporal patterns from historical tasks.
2 Related Work
2.1 Few-shot Image Classification
@@
The goal of few-shot image classification [25, 26, 27, 28] is to recognize unseen categories with limited annotated samples. Existing few-shot solutions can be broadly categorized into three groups: i) Augmentation-based methods typically generate auxiliary data samples by feature augmentation [29, 30, 31] or generative models [32, 33], to increase the diversity of the support set; ii) Optimization-based methods [34, 35, 36, 37, 38] learn good initial model parameters, such that updating the initial parameters via a few gradient steps could adapt well to new tasks; iii) Metric-based methods [39, 40, 41, 42] (e.g., ProtoNet [43] and RelationNet [44]) first learn a common feature space for all classes, and then employ pre-defined or learned distance metrics to compare query and support samples.
@@
Our work shares a similar spirit of metric-based methods [43, 45], whereas we address more complex and challenging few-shot action recognition that requires long-range relation modeling between spatio-temporal locations. With respect to this, we further jointly learn inter-video and inter-task relationships to learn distinct temporal features. By unifying such three types of relation modeling under our framework, we can learn task-specific temporal cues for each video from a holistic view.
2.2 Few-shot Action Recognition
@@
Few-shot Action Recognition (FSAR) is gaining much attention due to its practical value in reducing manual annotations at a large scale required by traditional action recognition. Compared with few-shot image classification, FSAR is more challenging due to the rich spatial-temporal information in videos. Towards FSAR, most existing solutions [46, 47, 48, 49] belong to the meta-learning paradigm [43, 45], wherein the model learns discriminative temporal features from base action categories, and generalizes learned temporal representations on novel action categories. Based on the meta-learning policy, a main group of efforts focus on temporal representation learning. Early works [9, 50] directly aggregate frame-level features to learn a global representation for each video. Though straightforward, these methods neglect complex temporal cues within one video, leading to suboptimal performance. To address this limitation, recent methods attempt to model temporal relations by, e.g., learning long-range temporal dependencies [14, 13], detailed cross-frame patch-level interactions [51], exploiting low-level spatial features via feature architecture search [16], or multi-modal feature fusion [52, 17, 53, 54]. Apart from investigating video feature representation, another line of work turns to designing various feature matching strategies, e.g., frame-level feature alignment [13], tuple-level feature alignment [19], and even multi-level feature alignment [21, 47, 55]. For example, OTAM [10] performs frame-level feature alignment between videos with the DTW algorithm, which takes the temporal ordering characteristics of videos into account. HyRSM [13] designs a novel bidirectional Mean Hausdorff metric to obtain temporal matching scores between support and query videos from the set matching perspective.
@@
Despite promising results have been achieved, they typically learn temporal representations independently for each video, ignoring the relation modeling between video samples and tasks. As a result, they fail to gain a holistic semantic pattern understanding from multi-level perspectives. Though a few works (e.g., GgHM [15] and HyRSM [13]) attempt to learn task-specific embeddings by considering inter-task relationships, they only focus on the correlations between samples at the video level, overlooking fine-grained temporal interactions among videos. In contrast, our HR2G-shot i) learns inter-video relationships at fine-grained frame level, and ii) gather temporal knowledge from historical tasks to achieve inter-task knowledge transfer, thereby capturing task-specific temporal features for each video and ensuring a sufficient understanding of holistic semantic patterns for each task.
3 Method
3.1 Problem Formulation
@@
The objective of few-shot action recognition (FSAR) is to classify novel action categories with limited samples per class. Under the few-shot setting, the model generalizes learned knowledge on the training set $\mathcal{D}_{base}$ to novel testing set $\mathcal{D}_{novel}$, where base classes $\mathcal{C}_{base}$ and novel classes $\mathcal{C}_{novel}$ are disjoint. Consistent with previous works [10, 13, 56], we formulate FSAR problem with episodic training and testing. Specifically, each $N$-way $K$-shot episode task involves sampled $K$ labeled videos from each of $N$ different action classes as the support set, and a portion of the remaining samples from $N$ classes as the query set. During inference, we randomly sample multiple episodic tasks from $\mathcal{D}_{novel}$ and present the average results over these tasks, in order to comprehensively evaluate the performance of the few-shot model.
[…]
Figure 2: The overview of HR2G-shot. (a) HR2G-shot unifies three types of relation modeling (i.e., inter-frame, inter-video, and inter-task) to learn discriminative temporal features. (b) Inter-video Semantic Correlation (ISC) conducts fine-grained cross-video interactions to learn inter-video relationships. (c) To explore inter-task relationships, we retrieve and aggregate temporal knowledge from the bank, which maintains diverse temporal patterns from historical tasks.
3.2 Algorithm Overview
@@
We introduce HR2G-shot, which hierarchically models multi-level relations (i.e., inter-frame, inter-video, and inter-task) to capture task-specific temporal patterns from a holistic view for FSAR. The overall architecture of HR2G-shot is shown in Fig. 2. Given an $N$-way $K$-shot episode task with $NK$ support videos and $L$ query videos, we employ CLIP visual encoder [57] to get frame-level features for support and query videos, i.e., $\bm{X}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{X}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$, where $T$ is the length of sampled frames in each video. Note that we adopt the Parameter-Efficient Fine-Tuning (PEFT) strategy to fine-tune the visual encoder with minimal trainable parameters as in [58, 59, 60]. Next, following previous works [61], we concatenate frame-level features and corresponding textual features along the temporal dimension and then feed them into the temporal Transformer [61, 18] (i.e., Inter-frame Temporal Modeling). In this way, we perform inter-frame interactions and obtain enhanced support and query features. To compensate for frame-level relation modeling from other semantic perspectives, we further design two feasible modules: i) Inter-video Semantic Correlation (ISC) (§3.3) learns inter-video temporal relationships within each task by fine-grained cross-video interactions; ii) Inter-task Knowledge Transfer (IKT) (§3.4) aggregates useful temporal knowledge from historical tasks to learn inter-task relationships. Finally, the obtained task-specific support and query features, i.e., $\bar{\bm{F}}_{\mathrm{s}}$ and $\bar{\bm{F}}_{\mathrm{q}}$, are fed into temporal matching metrics to get the class predictions.
3.3 Inter-video Semantic Correlation